What is the pOH of a solution that has a hydroxide ion concentration of 2.82 x 10-8 M?


1
Expert's answer
2021-08-20T01:51:17-0400

pH + pOH = 14

pOH = -lg [OH-] = -lg [2.82 x 10-8] = 7.55

pH = 14 - 7.55 = 6.45